You have a bag of marbles. There are twice as many blue marbles as red marbles. There are 4 less green marbles than red marbles. If there are 28 marbles in the bag, how many of each marble is in the bag? USE THE FIVE STEP PROBLEM SOLVING APPROACH.
PLEASE help me, I have been stuck on this problem for so long, and I really need to get it done, but I don't even know where to begin.
Thanks! :)

Answer by edjones(7569)

b=2r
g=r-4
r+g+b=28
r+r-4+2r=28 substitute r for g & b.
4r-4=28
4r=32
r=8
b=16
g=4
.
Ed